Grade 11General Physics

100 ml of 50% w/w HCL (aq) solutions.Density 1.46 g/ml is mixed with 100 ml of water calculate Molarity.

Anand
8 Years agoGrade 11

2 Answers

Norman Manchu
ApprovedApproved Tutor Answer8 Years ago
Molarity of mixture of two or more components is given by-​
Mtotal=\frac{M1V1+M2V2}{V1+V2}
 
For calculating M1 which is the molarity of HCl we can use the formula-
Molarity=\frac{10x\rho}{molecular weight}
Where x is percent weight and rho indicates density
M1=2x10^-1
Molarity of water is calculated by n(water)/V Thus-M2=5x10^-2
The total molarity can hence be calculated and the final answer would be 125x10^-3Molar
Kapil Hitesh Gupta
3 Years ago
This problem is a classical problem of dilution(as it involves decrease in conc. of sol. or change in volume due to only increase in volume of solvent).
 
So, there is a general and intuitive concept in dilution that Moles of solute before dilution = Moles of solute before dilution
This is so as in dilution we have two ways to dilute the solution:
1. Decrease the solute
2. Increase the volume of solution/solvent(while solute is const.)
In 99.99% cases we go with 2nd method as its practically way more easier than 1
 
Here's how we approach the answer to problem
 Before this, formulas you need to know M= nsolute/Volume of solution(L)
 
i.e., n =M×Vol.
 
According to the concept M1V1=M2V2
M1= Molarity of solution before
M2= Molarity of solution after
V1= Volume of solution before
V2= Volume of solution after
 
Mass of solution = 100 × 1.46 = 146 g (m= density×volume.)
 
As 50% (W/W) solution 73 g HCl and 73 g H2O
So, n solute = 73/36.5 = 2 
M1 = 2/0.1L=20M and V1=100ml, V2=200(100ml sol.+ 100ml added water
 
M1V1=M2V2
 
20×100ml = M2×200ml
M210M
Therfore the molariy of solution after dilution with 100ml water is 10 molar/M
 
NOTE: Be careful with the answers you see as the other answer provided is wrong(Though I'm not an IITian but I can clearly tell right and wrong)